OpenStudy (anonymous):

can some one plz tell me how i would do this? simplify the expression 5^7 × 5^11

OpenStudy (kinggeorge):

In an equation with exponents such as these, you can just add the exponents! So \[2^4\times 2^7=2^{4+7}\]In general, \[a^b\times a^c=a^{b+c}\]With this information, could you tell me the answer to your question?

OpenStudy (anonymous):

5^17?

OpenStudy (kinggeorge):

Almost, it would actually be 5^18 Since you have \(11+7=18\) instead of \(10+7=17\)

OpenStudy (anonymous):

oh...ok THX

OpenStudy (kinggeorge):

You're welcome.